OpenStudy (anonymous):

9 radical 7 divided by radical 28?? Help!?? :)

OpenStudy (anonymous):

start off with your equation: \[9\sqrt{7}\div \sqrt{28}\] now radical 28 is the same as radical 4(7): \[9\sqrt{7}\div \sqrt{4(7)}\] The square root of 4 is 2, seven has no square root: \[9\sqrt{7}\div2\sqrt{7}\] Now your root 7's will cancel out: \[9\div2\] and 9/2 is 4.5: 4.5 is your answer :D